OpenStudy (anonymous):

Evaluate the algebraic expression for x = -2 -5x2 - 3x my answer is -26

pooja195 (pooja195):

−5(−2)2−(3)(−2) =(−5)(4)−(3)(−2) =−20−(3)(−2) =−20−(−6) 2 negatives next to each other make a positive. -20+6=?

OpenStudy (anonymous):

-14?

pooja195 (pooja195):

Correct